OpenStudy (anonymous):

Solve the following equation: x^3+6x^2+15x+10=0

OpenStudy (anonymous):

Real: -1 Imaginary: (-5+i*sqrt(15))/2 and (-5-i*sqrt(15))/2

OpenStudy (anonymous):

The expression factors into the following:\[(x+1) \left(x^2+5 x+10\right)=0 \]x=-1 and:\[\left\{\left\{x\to \frac{1}{2} \left(-5-i \sqrt{15}\right)\right\},\left\{x\to \frac{1}{2} \left(-5+i \sqrt{15}\right)\right\}\right\} \]